OpenStudy (anonymous):

9^x-7.3^x+10=0,what is the minimum value wich made this equation correct?given(log2=0.3,log3=0.48)

OpenStudy (anonymous):

3^x=u get it!

OpenStudy (anonymous):

u²-7.u+10=0

OpenStudy (experimentx):

Hmm ... was that 3^x ??not 7.3^x ??

OpenStudy (anonymous):

yes my bad

OpenStudy (experimentx):

i was wondering it simply was untrue ... lol

OpenStudy (anonymous):

3^x=2

OpenStudy (experimentx):

x = log 2 / log 3

OpenStudy (anonymous):

x=.3/.48

OpenStudy (experimentx):

I bet the minimum value that would make this equation true would be http://www.wolframalpha.com/input/?i=x^2+-+6.3x+%2B+10+%3D+0+

OpenStudy (anonymous):

the minimum is 0.625

OpenStudy (experimentx):

Hmmm ... is it asking for two values of x??

OpenStudy (anonymous):

no only the minimum

OpenStudy (experimentx):

Oh ... i guess it is solved as like you did then